OpenStudy (angel_kitty12):

URGENT: I need help ASAP (4^1/2•2^1/3)^-8 Can someone please help I'm so lost thank you :)

OpenStudy (anonymous):

does it help to know that \(4^{\frac{1}{2}}=\sqrt4=2\)?

OpenStudy (angel_kitty12):

Yes well I'm a bit confused. Do I solve the problem in parenthesis first then multiply the exponent on the outside?

OpenStudy (anonymous):

we got \[(2\times 2^{\frac{1}{3}})^{-8}\] right ?

OpenStudy (angel_kitty12):

But then again, the inside would give me 2• 2• 3 cube root 2 which confuses me now.

OpenStudy (angel_kitty12):

I understand that so far yes

OpenStudy (anonymous):

add the exponents inside get \[\left(2^{\frac{4}{3}}\right)^{-8}\]

OpenStudy (anonymous):

on account of \(1+\frac{1}{3}=\frac{4}{3}\)

OpenStudy (angel_kitty12):

Wait how did you get 4?

OpenStudy (anonymous):

then multiply the exponents

OpenStudy (angel_kitty12):

How would you add the exponents 1+1/3 and get 4/3?

OpenStudy (anonymous):

\[\huge 2\times 2^{\frac{1}{3}}=2^1\times 2^{\frac{1}{3}}=2^{1+\frac{1}{3}}\]

OpenStudy (anonymous):

and \[1+\frac{1}{3}=\frac{4}{3}\]

OpenStudy (anonymous):

are you asking why \(1+\frac{1}{3}=\frac{4}{3}\)?

OpenStudy (anonymous):

cause i can explain that if you like

OpenStudy (angel_kitty12):

But wouldn't it be 1 1/3 ?

OpenStudy (anonymous):

yes same number

OpenStudy (angel_kitty12):

Yes I don't understand how you got 4/3 when adding 1+ 1/3

OpenStudy (anonymous):

do you know how to turn a mixed number in to an "improper fraction"?

OpenStudy (anonymous):

\[\huge 1\tfrac{1}{3}=\frac{3\times 1+1}{3}=\frac{4}{3}\]

OpenStudy (angel_kitty12):

Oh wait it's the same thing when divided into decimal format okay I understand

OpenStudy (angel_kitty12):

I understand that part okay

OpenStudy (anonymous):

nothing to do with decimals, just adding a fraction and a whole number

OpenStudy (anonymous):

finally multiply \(\frac{4}{3}\times (-8)\)

OpenStudy (angel_kitty12):

Then 4/3 multiply it by -8 to get -4/3 so it would be 4^-4/3 correct?

OpenStudy (anonymous):

\[\frac{4}{3}\times (-8)\neq -\frac{4}{3}\]

OpenStudy (anonymous):

\[4\times 8=32\] so \[\frac{4}{3}\times (-8)=-\frac{32}{3}\]

OpenStudy (angel_kitty12):

That's a large number. Okay I understand that

OpenStudy (anonymous):

it is not large first off it is negative second it is \(-10\tfrac{2}{3}\)

OpenStudy (anonymous):

in any case your answer is \[\huge 2^{-\frac{32}{3}}\]

OpenStudy (angel_kitty12):

Oh.

OpenStudy (angel_kitty12):

So the correct answer would be -10^2/3 or 2^-32/3 or are they both considered the same?

OpenStudy (anonymous):

answer would not be \[-10^{\frac{ 2 }{ 3 }}\] satellite was merely explaining that \[-\frac{ 32 }{ 3 }=-10\frac{ 2 }{ 3 }\] The correct answer would be \[2^-{\frac{ 32 }{ 2 }} or \frac{ 1 }{( \sqrt{2}^{32}) }\]

OpenStudy (angel_kitty12):

Thank you very much I understand now :)